OpenStudy (asapolous):

Write an algebraic equation for the following problem and then solve it. When Angela and Walker first started working for the​ supermarket, their weekly salaries totaled ​$650. Now during the last 25 years Walker has seen his weekly salary triple. Angela has seen her weekly salary become four times larger. Together their weekly salaries now total ​$2370. How much did they each make 25 years​ ago?

ganeshie8 (ganeshie8):

So what are we trying to find here ?

OpenStudy (asapolous):

we're trying to find and equation so something = 2370

OpenStudy (asapolous):

once there is an equation, i can solve it

ganeshie8 (ganeshie8):

Yes, we want to find the Old salaries of Angela and Walker by setting up an equation.

ganeshie8 (ganeshie8):

Let : A = old salary of Angela W = old salary of Walker

ganeshie8 (ganeshie8):

Sum of their old salaries was 650. Can you write an equation for this using above variables ?

OpenStudy (asapolous):

i cant

ganeshie8 (ganeshie8):

That's okay. You'd find it easier once I showed you how to setup the equation.

ganeshie8 (ganeshie8):

We have defined A and W for their past salaries. Since we know that their old salaries add up to 650 : A + W = 650

ganeshie8 (ganeshie8):

That's one equation. We need to setup another equation like that for the sum of their present salaries

ganeshie8 (ganeshie8):

What does their present salaries add up to ?

OpenStudy (asapolous):

2370

OpenStudy (asapolous):

3w+4a=2370

OpenStudy (asapolous):

?

ganeshie8 (ganeshie8):

Awesome!

ganeshie8 (ganeshie8):

So we have two equations : a + w = 650 3w+4a=2370

ganeshie8 (ganeshie8):

You should be able to solve them ?

OpenStudy (asapolous):

i cant use a and w as the variables, it tells me to type the expression as x as the variable

ganeshie8 (ganeshie8):

x for Angela or Walker ?

OpenStudy (asapolous):

so its, 3x+4x=2370

OpenStudy (asapolous):

but that would imply they make the same

ganeshie8 (ganeshie8):

No, can you take a screenshot of the question and attach here ?

ganeshie8 (ganeshie8):

Try this : 3x + 4(650-x) = 2370

ganeshie8 (ganeshie8):

Try entering this : 3x + 4*(650-x) = 2370